Baseballs are packed one dozen per box. There are 208 baseballs to be packed. How many more baseballs will be needed to fill the last partially filled box. Please show me how to express in an equation. Thank you.
Answer by ankor@dixie-net.com(22740)

Baseballs are packed one dozen per box. There are 208 baseballs to be packed. How many more baseballs will be needed to fill the last partially filled box. Please show me how to express in an equation. Thank you.
:
Let r = no. of balls required to fill the last box
:
r = 12 balls - (208 - (integer value of (208/12) * 12)
:
r = 12 - (208 - (int(208/12) * 12))
:
r = 12 - (208 - (17 * 12))
:
r = 12 - (208 - 204)
:
r = 12 - 4
:
r = 8 balls to fill the last box